# A Full AGN Feedback Prescription for Numerical Models: Negative, Positive and Hot Gas-Ejection Modes

### v1

@@ 5 Conclusions
Building upon the semianalytic model introduced in Contini et al. (2024b) and later updated in Contini et al. (2024a), we developed a new version named FEGA25, which introduces a comprehensive revision of AGN feedback, now split into three distinct modes: negative, positive, and hot gas ejection feedback. The first mode is the classical mechanism used in semianalytics and numerical simulations, which prevents hot gas from overcooling, thereby inhibiting star formation. The second is a more physically motivated version of the mechanism presented in Contini et al. (2024b), accounting for star formation induced by AGN activity. The third mode represents the primary novelty of this work: a mechanism capable of expelling hot gas beyond the virial radius of the halo.
FEGA25 successfully reproduces the galaxy properties and scaling relations explored in Contini et al. (2024b), while delivering significant improvements in the description of the main sequence of star-forming galaxies and the red fraction as a function of stellar mass. The model predicts a positive feedback mode that is more active at high redshifts, gradually fading over time as the negative mode becomes dominant. On average, the positive mode is more effective in galaxies with increasing stellar mass. Thanks to the introduction of the third feedback mode—the hot gas ejection mode—the fraction of hot gas in halos, from Milky Way-like scales to large clusters, aligns more closely with observational data, marking a significant advancement in the predictive power of our semianalytic model.
In conclusion, echoing the statement of Popesso et al. (2024), an additional physical mechanism capable of expelling gas beyond the halo while preserving the cold gas and stellar components is essential for numerical methods. The hot gas ejection mode introduced in FEGA25 appears to be a promising solution, effectively reducing the hot gas content within halos and improving agreement with observations. While there is still room for refinement, this mechanism lays a solid foundation, and we aim to further explore its implications in an upcoming paper.
@@ Acknowledgements
E.C. and S.K.Y. acknowledge support from the Korean National Research Foundation (2020R1A2C3003769). E.C. and S.J. acknowledge support from the Korean National Research Foundation (RS-2023-00241934). All the authors are supported by the Korean National Research Foundation (RS-2022-NR070872). J.R. was supported by the KASI-Yonsei Postdoctoral Fellowship and by the Korea Astronomy and Space Science Institute under the R&D program (Project No. 2023-1-830-00), supervised by the Ministry of Science and ICT.

### v2

@@ 4 Conclusions
Building upon the semianalytic model introduced in Contini et al. (2024b) and later updated in Contini et al. (2024a), we developed a new version named FEGA25, which introduces a comprehensive revision of AGN feedback, now split into three distinct modes: negative, positive, and hot gas ejection feedback. The first mode is the classical mechanism used in semianalytics and numerical simulations, which prevents hot gas from overcooling, thereby inhibiting star formation. The second is a more physically-motivated version of the mechanism presented in Contini et al. (2024b), accounting for star formation induced by AGN activity. The third mode represents the primary novelty of this work: a mechanism capable of expelling hot gas beyond the virial radius of the halo.
FEGA25 successfully reproduces the galaxy properties and scaling relations explored in Contini et al. (2024b), while delivering significant improvements in the description of the main sequence of star-forming galaxies and the passive fraction as a function of stellar mass. The model suggests a positive feedback mode that is more active at high redshifts, gradually fading over time as the negative mode becomes dominant. On average, the positive mode is more effective in galaxies with increasing stellar mass. Thanks to the introduction of the third feedback mode—the hot gas ejection mode—the fraction of hot gas in halos, from Milky Way-like scales to large clusters, aligns more closely with observational data, marking a significant advancement in the predictive power of our semianalytic model.
In conclusion, echoing the statement of Popesso et al. (2024), an additional physical mechanism capable of expelling gas beyond the halo while preserving the cold gas and stellar components is essential for numerical methods. The hot gas ejection mode introduced in FEGA25 appears to be a promising solution, effectively reducing the hot gas content within halos and improving agreement with observations. While there is still room for refinement, this mechanism provides a strong foundation. We plan to further explore its implications in an upcoming paper (Contini et al., in prep.), where we will examine the evolution of the baryon and hot gas fractions with redshift and disentangle the contributions of supernova and AGN feedback in expelling gas beyond the virial radius.
@@ Acknowledgements
E.C. and S.K.Y. acknowledge support from the Korean National Research Foundation (RS-2025-00514475). E.C. and S.J. acknowledge support from the Korean National Research Foundation (RS-2023-00241934). All the authors are supported by the Korean National Research Foundation (RS-2022-NR070872). J.R. was supported by the KASI-Yonsei Postdoctoral Fellowship and by the Korea Astronomy and Space Science Institute under the R&D program (Project No. 2023-1-830-00), supervised by the Ministry of Science and ICT.